Solve by addition
x - 2y = -1
-x + 5 y = 4

Answer by checkley75(3666)

x - 2y = -1
-x + 5 y = 4
-----------------------
3y=3
y=3/3
y=1 ans
x-2*1=-1
x-2=-1
x=-1+2
x=1 ans.
Proof:
-1+5*1=4
-1+5=4
4=4